# Don’t flatten, tokenize! Unlocking the key to SoftMoE’s efficacy in deep RL

### v1

@@ Abstract
Figure 1: Comparison of IQM (Agarwal et al., 2021) for Rainbow (Hessel et al., 2018), using the Impala architecture (Espeholt et al., 2018), with penultimate layer scaled, and with SoftMoE with varying numbers of experts, and a single, scaled, expert. IQM scores computed over 200M environment steps across 20 games, with 5 independent runs each. Error bars represent 95% stratified bootstrap confidence intervals. A single scaled expert matches the performance of multiple experts.
@@ 1 Introduction
Deep networks have been central to many of the successes of reinforcement learning (RL) since their initial use in DQN (Mnih et al., 2015a), which has served as the blueprint (in particular with regards to network architecture) for most of the followup works. However, recent works have highlighted the need to better understand the network architectures used, and question whether alternate choices can be more effective (Graesser et al., 2022; Sokar et al., 2022, 2023; Obando Ceron et al., 2024). Indeed, Obando Ceron* et al. (2024) and Willi* et al. (2024) demonstrated that the use of Mixtures-of-Experts (MoEs) (Shazeer et al., 2017), and in particular soft mixtures-of-experts (SoftMoEs) (Puigcerver et al., 2024), yields models that are more parameter efficient. Obando Ceron* et al. (2024) hypothesized that the underlying cause of the effectiveness of SoftMoEs was in the induced structured sparsity in the network architecture. While some of their analyses did seem to support this, the fact that performance gains were observed even with a single expert seems to run counter to this hypothesis.
@@ 2 Online Deep Reinforcement Learning
In online reinforcement learning, an agent interacts with an environment and adjusts its behaviour based on the reinforcement (in the form of rewards or costs) it receives from the environment. At timestep $t$, the agent inhabits a state $x_{t}\in\mathcal{X}$ (where $\mathcal{X}$ is the set of all possible states) and selects an action $a_{t}\in\mathcal{A}$ (where $\mathcal{A}$ is the set of all possible actions) according to its behaviour policy $\pi:\mathcal{X}\rightarrow\mathfrak{P}(\mathcal{A})$; the environment’s dynamics yield a new state $x_{t+1}\in\mathcal{X}$ and a numerical reward $r_{t}\in\mathbb{R}$. The goal of an RL agent is to find a policy $\pi^{*}$ that maximizes $V^{\pi}(x):=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}|x_{0}=x,a_{t}\sim\pi(x_{t})\right]$, where $\gamma\in[0,1)$ is a discount factor. In value-based RL, an agent maintains an estimate of $Q^{\pi}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, defined as $Q^{\pi}(x,a)=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}|x_{0}=x,a_{0}=a,a_{t}\sim\pi(x_{t})\right]$; from any state $x\in\mathcal{X}$, a policy $\pi$ can then be improved via $\pi(x)=\arg\max_{a\in\mathcal{A}}Q^{\pi}(x,a)$ (Sutton and Barto, 1998).
@@ 3 Mixtures of Experts
While there have been a number of architectural variations proposed, in this work we focus on Mixtures of Experts as used by Obando Ceron* et al. (2024). Mixtures of experts (MoEs) were originally proposed as a network architecture for supervised learning tasks using conditional computation, which enabled training very large networks (Shazeer et al., 2017; Fedus et al., 2022). This was achieved by including MoE modules which route a set of input tokens to a set of $n$ sub-networks, or experts; the output of these experts are then combined and fed through the rest of the outer network. The routing/gating mechanism and the expert sub-networks are all trainable. In top-$k$ routing, we consider two methods for pairing tokens and experts. In token choice routing, each token gets an assignation to $k$ experts, each of which has $p$ slots; this may result in certain tokens being dropped entirely by the MoE module and, more critically, certain experts may receive no tokens (Gale et al., 2023). To overcome having load-imbalanced experts, Zhou et al. (2022) propose expert choice routing, where each expert chooses $p$ tokens for its $p$ slots, which results in better expert load-balancing. See the bottom row of Figure 2 for an illustration of this. In most of our analyses, we focus on expert choice routing as it demonstrates higher performance than token choice routing (see Appendix B.1 for full empirical comparison between the two across varying number of experts).
@@
In order to maintain the shape of the input tokens (as originally done with MoE architectures), Obando Ceron* et al. add a projection layer after the “combine” step ($\Phi_{C}$ in Figure 2). Another important detail is the proposed tokenization scheme. MoEs are most commonly used in NLP transformer architectures where tokens are clearly defined (typically as encodings of words), but this tokenization is not something readily available for online RL agents. Thus, Obando Ceron* et al. proposed a number of mechanisms for tokenizing the output of the encoder.
@@ 4.2 Analysis of SoftMoE Components
At a high level, SoftMoE modules for online RL agents consist of (i) the use of a learnable tensor $\Phi$ used to obtain the dispatch ($\Phi_{D}$) and combine ($\Phi_{C}$) weights; (ii) processing $p$ input slots per expert; (iii) the architectural dimensions (network depth and width); (iv) the use of $n$ experts; and (v) the tokenization of the encoder output. The middle row of Figure 2 provides an illustration of these. In this section we discuss the first four with empirical evaluation demonstrated in Figures 1 and 4, and devote a separate section for tokenization, which turns out to be crucial for performance.
Figure 4: Understanding the impact of SoftMoE components. Using Rainbow with the Impala architecture as our base, we investigate key aspects of SoftMoE-4: (i) combining tokens, (ii) expert specialization, and (iii) adjusting architectural dimensions. SoftMoE does not appear to suffer from the performance degradation observed in the baseline, even when increasing its depth.
@@
Obando Ceron* et al. scaled up the penultimate layer of the baseline model with the intent of matching the parameters of the SoftMoE networks, thereby demonstrating that naïvely scaling width degrades performance. While they demonstrated that a performance drop is not observed when scaling down the dimensionality of each expert, they did not investigate what occurs when scaling up the experts dimensionality. This experiment makes sense if comparing the baseline’s penultimate layer dimensionality with that of each expert, rather than with the combined experts dimensionality. Scaling up the width of experts in SoftMoE does not lead to the same performance degradation observed in the scaled baseline model as shown in the first group [iii] of Figure 4.
@@
Figure 5: Evaluating impact of tokenizing Rainbow-lite with CNN (top) and Impala (bottom) architectures, using sum/mean over convolutional features. Tokenization with a scaled representation can yield significant performance gains.
@@ 5 Don’t flatten, tokenize!
Our first experiment in this series simply adds tokenization to the baseline architecture. Specifically, given the 3-dimensional encoder output with shape $[h,w,d]$, we replace the standard flattening (to a vector of dimension $h*w*d$) with a reshape to a matrix of shape $[h*w,d]$; then, prior to the final linear layer, we either sum over the first dimension or take their average. Since this operation is not immediately applicable to Rainbow (due to Noisy networks), we explored this experiment with Rainbow-lite, a version of Rainbow in Dopamine which only includes C51 (Bellemare et al., 2017), multi-step returns, and prioritized experience replay. Figure 5 demonstrates that this simple change can yield statistically significant improvements for Rainbow-lite, providing strong evidence that tokenization plays a major role in the efficacy of SoftMoE.
@@
Figure 6: Combined tokenization is a key driver of SoftMoE’s efficacy. SoftMoE with a single scaled expert has comparable performance with using multiple experts. Even non-combined tokenization (e.g. ExpertChoice) with a single expert achieves similar performance.
@@
We further investigate the impact of combined tokenization in the single expert setting. To this end, we study SoftMoE-1, with scaled expert dimensionality as before. As Figure 6 shows, using combined tokens with a single expert brings the performance even closer to the multiple expert setting. This further confirms that (combined) tokenization is the key driver of SoftMoE’s efficacy.
@@ 5.1 Additional analyses
Figure 7: Comparison of different tokenization schemes for SoftMoE-1 with the standard and scaled penultimate layer. PerConv achieves the highest performance, suggesting that maintaining the spatial structure of the features plays a role on the observed benefits.
Figure 8: Evaluating the impact of fewer slots with a single scaled (by 4x) expert for both SoftMoE and ExpertChoice. Even with drastically fewer slots, SoftMoE-1 remains competitive.
@@
Puigcerver et al. (2024) highlighted that the time complexity of SoftMoE depends directly on the number of slots. As can be seen in Figure 8, using SoftMoE-1 with only 10% of the default number of slots, and even with a single slot, achieves a comparable performance to using the regular number of slots. On the other hand, the reduced performance of ExpertChoice-1 further confirms the gains added by combining tokens in SoftMoE.
Figure 9: (Left) Evaluating the performance on all the 60 games of Atari 2600. Consistent with our results SoftMoE-1 matches the performance of multiple experts over the full suite. (Right) CNN Encoder. The performance gain of SoftMoE-1 is present when training with the standard CNN encoder, revealing the importance of tokenization across architectures.
@@
Our results so far were run on the 20 games used by Obando Ceron* et al. (2024) and with the Impala ResNet encoder (Espeholt et al., 2018). To ensure our findings are not specific to these choices, we extend our evaluation to all 60 games in the ALE suite, and with the standard CNN encoder of Mnih et al. (2015b). Our findings, as illustrated in Figure 9, demonstrate consistent results across all games and various encoder architectures.
Figure 10: Evaluating on DQN (top) and DER (bottom). When combined with DQN, SoftMoE with a single scaled expert does not provide gains over the baseline. In contrast, when combined with DER, it achieves performance comparable to SoftMoE-4.
@@
We focused on Rainbow for the majority of our investigation due to the positive results demonstrated by Obando Ceron* et al. (2024), but it is important to also consider other agents. To this point, we evaluate the impact of a single scaled expert on DQN (Mnih et al., 2015b). Note that we replace the original Huber loss used in DQN with mean-squared error loss, as it has been shown to be more performant (Ceron and Castro, 2021), and coincides with what was used by Obando Ceron* et al.. and DER (Van Hasselt et al., 2019). Following the training regime used by Obando Ceron* et al. (2024), we train DER for 50M environment frames. In the DQN setting, SoftMoE-1 does not appear to offer significant improvements over the baseline, potentially due to the limited gains observed with SoftMoE-4 (Figure 10, top). Further investigation is needed to understand how the inherent instability of DQN might limit the potential benefits of SoftMoE. On the other hand, consistent with our finding in Rainbow, in DER, where SoftMoE-4 achieves a superior performance over the baseline, SoftMoE-1 is able to offer the same gains (Figure 10, bottom).
@@ 6 Calling all experts!
As a first step in this direction, we adopt existing techniques designed to improve network utilization: (1) periodic reset of parameters (Nikishin et al., 2022) and (2) using Shrink and Perturb (S&P) (Ash and Adams, 2020), as used by D’Oro et al. (2022) and Schwarzer* et al. (2023). We perform these experiments on Rainbow and DER with SoftMoE-4. We explore applying resets and S&P to all experts and to a subset of experts, where the experts to reset are those that have the highest fraction of dormant neurons (Sokar et al., 2023). For Rainbow, we reset every 20M environment steps, whereas for DER we reset every 200k environment steps.
Interestingly, the effectiveness of these techniques varied across different reinforcement learning agents. While weight reset and S&P did not improve expert utilization in the Rainbow agent, they led to enhanced performance in the sample-efficient DER agent (Figure 11). Within the DER agent, resetting a subset of experts proved more beneficial than resetting all, while applying S&P to all experts yielded the most significant performance gains. These findings highlight the need for further research into methods that can optimize expert utilization and improve performance in reinforcement learning with SoftMoE.
Figure 11: Evaluating the use of resets (top) and S&P (bottom) for improving expert utilization on Rainbow (left column) and DER (right column). All experiments run with SoftMoE-4. Resets and S&P seem to benefit DER, but degrade performance for Rainbow.

### v2

@@ Abstract
Figure 1: Comparison of IQM (Agarwal et al., 2021) for Rainbow (Hessel et al., 2018), using the Impala architecture (Espeholt et al., 2018), with penultimate layer scaled, and with SoftMoE with varying numbers of experts, and a single, scaled, expert. IQM scores computed over 200M environment steps across 20 games, with 5 independent runs each, higher is better. Error bars represent 95% stratified bootstrap confidence intervals. A single scaled expert matches the performance of multiple experts.
@@ 1 Introduction
Deep networks have been central to many of the successes of reinforcement learning (RL) since their initial use in DQN (Mnih et al., 2015a), which has served as the blueprint (in particular with regards to network architecture) for most of the followup works. However, recent works have highlighted the need to better understand the network architectures used, and question whether alternate choices can be more effective (Graesser et al., 2022; Sokar et al., 2022; 2023; Obando Ceron et al., 2024). Indeed, Obando Ceron* et al. (2024) and Willi* et al. (2024) demonstrated that the use of Mixtures-of-Experts (MoEs) (Shazeer et al., 2017), and in particular soft mixtures-of-experts (SoftMoEs) (Puigcerver et al., 2024), yields models that are more parameter efficient. Obando Ceron* et al. (2024) hypothesized that the underlying cause of the effectiveness of SoftMoEs was in the induced structured sparsity in the network architecture. While some of their analyses did seem to support this, the fact that performance gains were observed even with a single expert seems to run counter to this hypothesis.
@@ 2 Online Deep Reinforcement Learning
In online reinforcement learning, an agent interacts with an environment and adjusts its behaviour based on the reinforcement (in the form of rewards or costs) it receives from the environment. At timestep $t$, the agent inhabits a state $x_{t}\in\mathcal{X}$ (where $\mathcal{X}$ is the set of all possible states) and selects an action $a_{t}\in\mathcal{A}$ (where $\mathcal{A}$ is the set of all possible actions) according to its behaviour policy $\pi:\mathcal{X}\rightarrow\mathfrak{P}(\mathcal{A})$; the environment’s dynamics yield a new state $x_{t+1}\in\mathcal{X}$ and a numerical reward $r_{t}\in\mathbb{R}$. The goal of an RL agent is to find a policy $\pi^{*}$ that maximizes $V^{\pi}(x):=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}|x_{0}=x,a_{t}\sim\pi(x_{t})\right]$, where $\gamma\in[0,1)$ is a discount factor. In value-based RL, an agent maintains an estimate of $Q^{\pi}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$, defined as $Q^{\pi}(x,a)=\sum_{t=0}^{\infty}\left[\gamma^{t}r_{t}|x_{0}=x,a_{0}=a,a_{t}\sim\pi(x_{t})\right]$; from any state $x\in\mathcal{X}$, a policy $\pi$ can then be improved via $\pi(x)=\arg\max_{a\in\mathcal{A}}Q^{\pi}(x,a)$ (Sutton & Barto, 1998).
@@ 3 Mixtures of Experts
While there have been a number of architectural variations proposed, in this work we focus on Mixtures of Experts as used by Obando Ceron* et al. (2024). Mixtures of experts (MoEs) were originally proposed as a network architecture for supervised learning tasks using conditional computation, which enabled training very large networks (Shazeer et al., 2017; Fedus et al., 2022). This was achieved by including MoE modules which route a set of input tokens to a set of $n$ sub-networks, or experts; the output of these experts are then combined and fed through the rest of the outer network. The routing/gating mechanism and the expert sub-networks are all trainable. In top-$k$ routing, we consider two methods for pairing tokens and experts. In token choice routing, each token gets an assignation to $k$ experts, each of which has $p$ slots (the number of processed inputs by each expert); this may result in certain tokens being dropped entirely by the MoE module and, more critically, certain experts may receive no tokens (Gale et al., 2023). To overcome having load-imbalanced experts, Zhou et al. (2022) proposed expert choice routing, where each expert chooses $p$ tokens for its $p$ slots, which results in better expert load-balancing. See the bottom row of Figure 2 for an illustration of this. Unless otherwise specified, we use a value of $p$ equals to #$\frac{\#tokens}{\#experts}$. In most of our analyses, we focus on expert choice routing as it demonstrates higher performance than token choice routing (see Appendix B.1 for full empirical comparison between the two across varying number of experts).
@@
In order to maintain the shape of the input tokens (as originally done with MoE architectures), Obando Ceron* et al. (2024) add a projection layer after the “combine” step ($\Phi_{C}$ in Figure 2). Another important detail is the proposed tokenization scheme. MoEs are most commonly used in NLP transformer architectures where tokens are clearly defined (typically as encodings of words), but this tokenization is not something readily available for online RL agents. Thus, Obando Ceron* et al. (2024) proposed a number of mechanisms for tokenizing the output of the encoder.
@@ 4.2 Analysis of SoftMoE Components
At a high level, SoftMoE modules for online RL agents consist of (i) the use of a learnable tensor $\Phi$ used to obtain the dispatch ($\Phi_{D}$) and combine ($\Phi_{C}$) weights; (ii) processing $p$ input slots per expert; (iii) the architectural dimensions (network depth and width); (iv) the use of $n$ experts; and (v) the tokenization of the encoder output. The middle row of Figure 2 provides an illustration of these. In this section we discuss the first four with empirical evaluation demonstrated in Figures 1 and 4, and devote a separate section for tokenization, which turns out to be crucial for performance. The same analysis was performed on DER, yielding the same results, and can be found in Appendix B.3.
Figure 4: Understanding the impact of SoftMoE components. Using Rainbow with the Impala architecture as our base, we investigate key aspects of SoftMoE-4: (i) combining tokens, (ii) expert specialization, and (iii) adjusting architectural dimensions. Reporting IQM (Agarwal et al., 2021), where higher is better. SoftMoE does not appear to suffer from the performance degradation observed in the baseline, even when increasing its width.
@@
Obando Ceron* et al. (2024) scaled up the penultimate layer of the baseline model with the intent of matching the parameters of the SoftMoE networks, thereby demonstrating that naïvely scaling width degrades performance. While they demonstrated that a performance drop is not observed when scaling down the dimensionality of each expert, they did not investigate what occurs when scaling up the experts dimensionality. This experiment makes sense if comparing the baseline’s penultimate layer dimensionality with that of each expert, rather than with the combined experts dimensionality. Scaling up the width of experts in SoftMoE does not lead to the same performance degradation observed in the scaled baseline model as shown in the first group [iii] of Figure 4.
@@
Figure 5: Evaluating impact of tokenizing Rainbow-lite with unscaled (top) and scaled (bottom) architectures, using sum/mean over convolutional features, and exploring both PerConv and PerFeat tokenization schemes. Reporting Median, IQM, Mean, and Optimaly Gap scores (Agarwal et al., 2021). Higher is better for all except Optimality Gap. Tokenization with a scaled representation can yield significant significant performance gains. Per-game results available in Appendix C.1.
@@ 5 Don’t flatten, tokenize!
Our first experiment simply adds tokenization to the baseline architecture. Specifically, given the 3-dimensional encoder output with shape $[h,w,d]$, we replace the standard flattening (to a vector of dimension $h*w*d$) with a reshape to a matrix of shape $[h*w,d]$ for PerConv tokenization or $[d,h*w]$ for PerFeat; then, prior to the final linear layer, we either sum over the first dimension or take their average. Since this operation is not immediately applicable to Rainbow (due to Noisy networks), we explored this experiment with Rainbow-lite, a version of Rainbow in Dopamine which only includes C51 (Bellemare et al., 2017), multi-step returns, and prioritized experience replay. Figure 5 demonstrates that this simple change can yield statistically significant improvements for Rainbow-lite, providing strong evidence that tokenization plays a major role in the successful scaling of DRL networks.
@@
Figure 6: Combined tokenization is a key driver of SoftMoE’s efficacy. Reporting IQM (Agarwal et al., 2021), where higher is better. SoftMoE with a single scaled expert has comparable performance with using multiple experts. Even non-combined tokenization (e.g. ExpertChoice) with a single expert achieves similar performance.
@@
We further investigate the impact of combined tokenization in the single expert setting. To this end, we study SoftMoE-1, with scaled expert dimensionality as before. As Figure 6 shows, using combined tokens with a single expert brings the performance even closer to the multiple expert setting. Similar results were observed in SoftMoE with 2 and 8 experts (See Appendix B.2 for more details). This further confirms that (combined) tokenization is the key driver of SoftMoE’s efficacy.
@@ 5.1 Additional analyses
Figure 7: Comparison of different tokenization schemes for SoftMoE-1 with the standard and scaled penultimate layer. Reporting IQM (Agarwal et al., 2021), where higher is better. PerConv achieves the highest performance, suggesting that maintaining the spatial structure of the features plays a role on the observed benefits.
@@
Figure 8: Evaluating the impact of fewer slots with a single scaled (by 4x) expert for both SoftMoE and ExpertChoice. Reporting IQM (Agarwal et al., 2021), where higher is better. Even with drastically fewer slots, SoftMoE-1 remains competitive.
@@
Puigcerver et al. (2024) highlighted that the time complexity of SoftMoE depends directly on the number of slots. As can be seen in Figure 8, using SoftMoE-1 with only 10% of the default number of slots (kept fixed throughout training), and even with a single slot, achieves a comparable performance to using the regular number of slots. On the other hand, the reduced performance of ExpertChoice-1 further confirms the gains added by combining tokens in SoftMoE.
@@
Our results so far were run on the 20 games used by Obando Ceron* et al. (2024) and with the Impala ResNet encoder (Espeholt et al., 2018). To ensure our findings are not specific to these choices, we extend our evaluation to all 60 games in the ALE suite, and with the standard CNN encoder of Mnih et al. (2015b). Our findings, as illustrated in Figure 9, demonstrate consistent results across all games and various encoder architectures. We further conducted extra experiments on Procgen (Cobbe et al., 2019), which are consistent with our findings (Figure 21 in Appendix C).
@@
We focused on Rainbow for the majority of our investigation due to the positive results demonstrated by Obando Ceron* et al. (2024), but it is important to also consider other agents. To this point, we evaluate the impact of a single scaled expert on DQN (Mnih et al., 2015b) and DER (Van Hasselt et al., 2019). Note that we replace the original Huber loss used in DQN with mean-squared error loss, as it has been shown to be more performant (Ceron & Castro, 2021), and coincides with what was used by Obando Ceron* et al. (2024). Following the training regime used by Obando Ceron* et al. (2024), we train DER for 50M environment frames. In the DQN setting, SoftMoE-1 does not appear to offer significant improvements over the baseline, potentially due to the limited gains observed with SoftMoE-4 (Figure 10, top). Further investigation is needed to understand how the inherent instability of DQN might limit the potential benefits of SoftMoE. On the other hand, consistent with our finding in Rainbow, in DER, where SoftMoE-4 achieves a superior performance over the baseline, SoftMoE-1 is able to offer the same gains (Figure 10, bottom). To further evaluate the generality of our claims, we conducted initial experiments with SAC (Haarnoja et al., 2018) on the CALE (Farebrother & Castro, 2024), which is a continuous action version of the original ALE. In Figure 22 in Appendix C it can be seen that tokenization provides an improvement over flattening.
Figure 9: (Left) Evaluating the performance on all of the 60 games of Atari 2600. Reporting IQM (Agarwal et al., 2021), where higher is better. Consistent with our results SoftMoE-1 matches the performance of multiple experts over the full suite. (Right) CNN Encoder. The performance gain of SoftMoE-1 is present when training with the standard CNN encoder, revealing the importance of tokenization across architectures.
@@ 6 Calling all experts!
Figure 10: Evaluating on DQN (top) and DER (bottom). Reporting Median, IQM, Mean, and Optimaly Gap scores (Agarwal et al., 2021). Higher is better for all except Optimality Gap. When combined with DQN, SoftMoE with a single scaled expert does not provide gains over the baseline. In contrast, when combined with DER, it achieves performance comparable to SoftMoE-4.
As a first step in this direction, we adopt existing techniques designed to improve network utilization: (1) periodic reset of parameters (Nikishin et al., 2022) and (2) Shrink and Perturb (S&P) (Ash & Adams, 2020), as used by D’Oro et al. (2022) and Schwarzer* et al. (2023). We perform these experiments on Rainbow and DER with SoftMoE-4. We explore applying resets and S&P to either all experts or a subset of experts, where the experts to reset are those that have the highest fraction of dormant neurons (Sokar et al., 2023). We reset every 20M and 200K environment steps for Rainbow and DER, respectively. More experimental details can be found in Appendix B.5.
Interestingly, the effectiveness of these techniques varied across different RL agents. While weight reset and S&P did not improve expert utilization in the Rainbow agent, they led to enhanced performance in the sample-efficient DER agent (Figure 11). Within the DER agent, resetting a subset of experts proved more beneficial than resetting all, while applying S&P to all experts yielded the most significant performance gains. These findings highlight the need for further research into methods that can optimize expert utilization and improve performance in RL with SoftMoE.
Figure 11: Evaluating the use of resets (top) and S&P (bottom) for improving expert utilization on Rainbow (left column) and DER (right column). Reporting IQM (Agarwal et al., 2021), where higher is better. All experiments run with SoftMoE-4. Resets and S&P seem to benefit DER, but degrade performance for Rainbow.